# Low-Rank Bandits via Tight Two-to-Infinity Singular Subspace Recovery

### v1

@@ C.3 Additional Lemmas
  Input: Samples $\left(i_{t},j_{t},r_{t}\right)_{t\in[T]}$, context distribution $\rho$, behavior policy $\pi^{b}$, target policy $\pi$, sample size $T_{1}$, regularization parameter $\tau$, confidence level $\delta$;
@@
  Output: $\hat{v}^{\pi}_{\text{\tt RS-PE}}=\sum_{i,j}w^{\pi}_{i,j}\phi_{i,j}^{T}\hat{\theta}$.
The general idea of the proof of Theorem 6.2 is to decompose the PE error of RS-PE as a sum of a term that depends on the misspecification error and a term incurred by the error between $\hat{\theta}$ and $\theta$. The analysis of the latter term, which is in fact the largest contributor to the overall error, is heavily inspired by the analysis in Appendix C.2. We first derive a high probability guarantee with the lower-order term that one would expect from the contextual linear bandit analysis, i.e., roughly $\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\sqrt{\frac{\log\left(1/\delta\right)}{T}}$. This bound is however not satisfactory in itself since it is a function of the randomness of the first phase of the algorithm. This issue can be resolved by leveraging the subspace recovery guarantees of Proposition 4.1 to show that $\|\phi_{\pi}\|_{2}$ converges towards $\|\psi_{\pi}\|_{2}$.
@@
Assume that $T_{1}=\lfloor\alpha T\rfloor$ for some $\alpha\in\left(0,1\right)$ and that the regularization parameter of $\widehat{\Lambda}_{\tau}$ verifies $\tau\leq\frac{r}{\left(m\wedge n\right)}\frac{\omega_{\min}}{\omega_{\max}}\log\left(16d/\delta\right)$. There exists an universal constant $C>0$ such that as long as $T\geq\frac{C}{\omega_{\min}mn}\mu^{4}\kappa^{2}r^{2}(m+n)\log^{3}\left(e(m+n)T/\delta\right)$ and $T\geq\frac{C}{\omega_{\min}}\log\left(16d/\delta\right)$, then with probability larger than $1-\delta$,
$\left|v^{\pi}-\hat{v}_{\text{\tt RS-PE}}\right|\leq\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\sqrt{\frac{2\log(16/\delta)}{\left(1-\alpha\right)T}}+\operatorname{poly}\left(m,n,\omega_{\min}^{-1},\mu,\kappa,r,\|M\|_{\max}\right)\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T}$
where $\operatorname{poly}\left(m,n,\omega_{\min}^{-1},\mu,\kappa,r,\|M\|_{\max}\right)\lesssim\max\left(1,\|M\|_{\max}^{2}\right)\mu^{8}\kappa^{11/2}r^{4}\frac{\left(m+n\right)^{2}\left(m\vee n\right)}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}.$
@@
$\leq\left|\frac{1}{T_{2}}\sum_{t=T_{1}+1}^{T}\phi_{\pi}^{\top}\left(\xi_{t}+\epsilon_{i_{t},j_{t}}\right)\widehat{\Lambda}_{\tau}^{-1}\phi_{i_{t},j_{t}}\right|+\frac{\tau}{T_{2}}\phi_{\pi}^{\top}\widehat{\Lambda}_{\tau}^{-1}\theta+\epsilon$
@@
$+\epsilon\left(1+\left|\frac{1}{T_{2}}\sum_{t=T_{1}+1}^{T}\phi_{\pi}^{\top}\widehat{\Lambda}_{\tau}^{-1}\phi_{i_{t},j_{t}}\right|\right).$
@@
$L_{3}=\epsilon\left(1+\left|\frac{1}{T_{2}}\sum_{t=T_{1}+1}^{T}\phi_{\pi}^{\top}\widehat{\Lambda}_{\tau}^{-1}\phi_{i_{t},j_{t}}\right|\right).$
As discussed before, we first perform the analysis of the error conditionally on the first phase of the algorithm. We will then control the misspecification error induced by the first phase using Proposition 5.1. Let $L:=\max_{i,j}\phi_{i,j}^{\top}\Lambda_{\phi}^{-1}\phi_{i,j}$, $L_{\theta}:=\max_{i,j}\left(\phi_{i,j}^{\top}\theta\right)^{2}$. Adapting the proof of Theorem C.7 yields the following lemma.
@@
$|L_{3}|\leq\epsilon\left(1+\sqrt{L}\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\left(1+4\sqrt{\frac{\log\left(16d/\delta\right)}{T_{2}}}\right)\right)$
with $E_{1}:=\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\sqrt{2\log(16/\delta)},E_{2}:=\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\left(LL_{\theta}+(1+4\sqrt{2d})\sqrt{L}\right)\log\left(32(d+1)/\delta\right),E_{3}:=\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\left(\frac{\sqrt{2}}{3}+4\sqrt{2d}+4\sqrt{L}L_{\theta}\right)L\log^{3/2}\left(32(d+1)/\delta\right),E_{4}:=\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}4\sqrt{2d}L^{3/2}\log^{2}\left(32(d+1)/\delta\right).$
@@
$\left|L_{2}\right|$ $\leq\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\left(\|\Lambda_{\phi}^{1/2}\Delta\Lambda_{\phi}^{1/2}\|_{{\textup{op}}}\|\Lambda_{\phi}^{-1/2}\hat{W}\|_{2}+\frac{\tau L_{\theta}}{\lambda_{\min}\left(\Lambda_{\phi}\right)T_{2}}\left(1+\|\Lambda^{1/2}\Delta\Lambda_{\phi}^{1/2}\|_{{\textup{op}}}\right)\right).$
for $\hat{W}=\frac{1}{T_{2}}\sum_{t=T_{1}+1}^{T_{2}}\xi_{t}\phi_{i_{t},j_{t}},$ The only difference is the additional term $L_{3}$, which can be controlled similarly to $L_{2}$. Indeed, by Cauchy-Schwartz and the definition of the operator norm,
@@
$\leq\frac{1}{T_{2}}\sum_{t=T_{1}+1}^{T}\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\sqrt{\phi_{i,j}^{\top}\Lambda_{\phi}^{-1}\phi_{i,j}}$
$+\frac{1}{T_{2}}\sum_{t=T_{1}+1}^{T}\|\Lambda_{\phi}^{-1/2}\phi_{\pi}\|_{2}\|\Lambda_{\phi}^{1/2}\Delta\Lambda_{\phi}^{1/2}\|_{2}\|\Lambda_{\phi}^{-1/2}\phi_{i_{t},j_{t}}\|_{2}$
$\leq\sqrt{L}\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\left(1+\|\Lambda_{\phi}^{1/2}\Delta\Lambda_{\phi}^{1/2}\|_{2}\right).$
Thus, $L_{3}\leq\epsilon\left(1+\sqrt{L}\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\left(1+\|\Lambda_{\phi}^{1/2}\Delta\Lambda_{\phi}^{1/2}\|_{2}\right)\right).$
By the same arguments as in the proof of Lemma B.6. from (Duan et al., 2020), if $T_{2}\geq 18L\log\left(8d/\delta\right)$ and $\tau\leq L\log(8d/\delta)\lambda_{\min}\left(\Lambda_{\phi}\right)$, then $\|\Lambda_{\phi}^{1/2}\Delta\Lambda_{\phi}^{1/2}\|_{2}\leq 4\sqrt{\frac{\log(8d/\delta)L}{T_{2}}}$ with probability larger than $41-\delta/$ 4.
We can now derive high probability bounds on $L_{1}$, $L_{2}$ and $L_{3}$ by adapting the contextual linear bandit analysis. The main difference is that here, the analysis is performed conditionally on the randomness of the first phase of the algorithm. One subtlety to note is that we need to use a version of Proposition C.9 that is conditional on the first $T_{1}$ samples because the choices of $\sigma$ and $\varepsilon$ in the analysis depend on $\phi$, which depends on the estimator $\widehat{M}$ built with the first $T_{1}$ samples. It is not an issue as it directly implies an unconditional high probability bound by taking the expectation.
Following the proof of Theorem C.7 thus yields the high probability bound in the statement as soon as $T_{2}\geq 18L\log\left(16d/\delta\right)$ and $\tau\leq L\log(16d/\delta)\lambda_{\min}\left(\Lambda_{\phi}\right)$. All that remains to be checked is that these inequalities are true as soon as $T_{2}\geq\frac{18}{\omega_{\min}}\log\left(16d/\delta\right)$ and $\tau\leq\frac{r}{\left(m\wedge n\right)}\frac{\omega_{\min}}{\omega_{\max}}\log\left(16d/\delta\right)$. By Lemma D.11 and the fact that $\max_{i,j}\|\phi_{i,j}\|^{2}_{2}\geq\max\left(\|U\|^{2}_{2\to\infty},\|V\|^{2}_{2\to\infty}\right)$,
$\frac{r}{\left(m\wedge n\right)\omega_{\max}}\leq\max_{i,j}\frac{\|\phi_{i,j}\|^{2}_{2}}{\lambda_{\max}\left(\Lambda_{\phi}\right)}\leq L\leq\max_{i,j}\frac{\|\phi_{i,j}\|^{2}_{2}}{\lambda_{\min}\left(\Lambda_{\phi}\right)}\leq\frac{1}{\omega_{\min}},$
@@
Note that in addition to $L\leq\frac{1}{\omega_{\min}}$, we also have $L_{\theta}\leq\max_{i,j}\|\phi_{i,j}\|^{2}_{2}\|\theta\|^{2}_{2}\leq mn\|M\|^{2}_{\max}$. This shows that the terms that depend on $E_{2}$, $E_{3}$ and $E_{4}$ are of lower order compared to the term that depends on $E_{1}$, and that they have a polynomial dependency in the problem parameters (note that $\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\leq\sqrt{L}$). We will now see that we can obtain a sharper control of the dependency in the problem parameters by leveraging high probability bounds on $L$ and $L_{\theta}$ instead of deterministic ones. These high probability bounds come freely with the control of the misspecification error.
@@
Let $\delta>0$. There exists an universal constant $c>0$ such that $T_{1}\geq\frac{c\mu^{4}\kappa^{2}r^{2}\left(m+n\right)}{\omega_{\min}mn}\log^{3}\left((m+n)T_{1}/\delta\right)$ ensures that there exists three universal constants $C_{0}$, $C_{1}$ and $C_{2}$ such that with probability larger than $1-\delta$, the following inequalities happen at the same time :
  • $\epsilon\leq\frac{C_{0}\|M\|_{\max}\mu^{6}\kappa^{4}r^{3}\left(m+n\right)\log^{3}\left((m+n)T_{1}/\delta\right)}{T_{1}\omega_{\min}\sqrt{mn}\left(m\wedge n\right)}$
  • $L_{\theta}\leq C_{1}\|M\|^{2}_{\max}\mu^{4}\kappa^{4}r^{2}\frac{m\vee n}{m\wedge n}$
  • $L\leq C_{2}\frac{\mu^{2}r\kappa\left(m+n-\mu^{2}r\right)}{\omega_{\min}mn}$.
@@
Proposition 5.1 corresponds to the first inequality. Note that this result only uses the high probability bounds on $\max\left(\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{2\to\infty},\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{2\to\infty}\right)$ of Proposition 4.1. This Proposition ensures that there is an universal constant $C$ such that the event
$A_{\delta}=\left\{\max\left(\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{2\to\infty},\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{2\to\infty}\right)\leq C\mu^{3}\kappa^{2}r^{3/2}\sqrt{\frac{m+n}{T_{1}mn\left(m\wedge n\right)\omega_{\min}}}\log^{3/2}\left(\frac{\left(m+n\right)T_{1}}{\delta}\right)\right\}$
holds with probability at least $1-\delta.$ Under $A_{\delta}$, we know from the proof of Proposition 5.1 that
@@
$\epsilon\leq\frac{C_{0}\|M\|_{\max}\mu^{6}\kappa^{4}r^{3}\left(m+n\right)\log^{3}\left((m+n)T_{1}/\delta\right)}{T_{1}\omega_{\min}\sqrt{mn}\left(m\wedge n\right)}$
@@
for some universal constant $C_{0}$. Let us now prove that both of the remaining inequalities are true under $A_{\delta}.$ Specifically, we will see that $L_{\theta}$ can be directly controlled by $\epsilon$, and that $L$ can be deterministically bounded by $\max\left(\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{2\to\infty},\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{2\to\infty}\right)$.
@@
Control of $L_{\theta}$. Recall that $M_{i,j}=\phi_{i,j}^{\top}\theta+\epsilon_{i,j}.$ Consequently, $\sqrt{L_{\theta}}=\displaystyle\max_{i,j}\left|\phi_{i,j}^{\top}\theta\right|\leq\|M\|_{\max}+\epsilon$. Since $T_{1}\geq\frac{c\mu^{4}\kappa^{2}r^{2}\left(m+n\right)}{\omega_{\min}mn}\log^{3}\left((m+n)T_{1}/\delta\right)$, under $A_{\delta}$ we have $\epsilon\leq\frac{C_{0}\|M\|_{\max}\mu^{2}\kappa^{2}r\sqrt{mn}}{c\left(m\wedge n\right)}$, which entails
$L_{\theta}\leq C_{1}\|M\|^{2}_{\max}\mu^{4}\kappa^{4}r^{2}\frac{m\vee n}{m\wedge n}$
for some universal constant $C_{1}.$
@@
Control of $L$. Recall that $L\leq\displaystyle\max_{i,j}\frac{\|\phi_{i,j}\|^{2}_{2}}{\lambda_{\min}\left(\Lambda_{\phi}\right)}\leq\max_{i,j}\frac{\|\phi_{i,j}\|^{2}_{2}}{\omega_{\min}}.$ Note that by Lemma D.12, $\max_{i,j}\|\phi_{i,j}\|^{2}_{2}\leq\mu^{2}r\frac{m+n-\mu^{2}r}{mn}+3\max\left(\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{2\to\infty},\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{2\to\infty}\right)$. Consequently,
$L\leq\frac{1}{\omega_{\min}}\left(\mu^{2}r\frac{m+n-\mu^{2}r}{mn}+3\max\left(\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{2\to\infty},\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{2\to\infty}\right)\right).$
Note that since $T_{1}\geq\frac{c\mu^{4}\kappa^{2}r^{2}\left(m+n\right)}{\omega_{\min}mn}\log^{3}\left(e(m+n)T_{1}/\delta\right)$, under $A_{\delta}$, we have
$\max\left(\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{2\to\infty},\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{2\to\infty}\right)\leq C\mu^{3}\kappa^{2}r^{3/2}\frac{1}{\sqrt{c\mu^{4}\kappa^{2}r^{2}\left(m\wedge n\right)}}=\frac{C\mu\kappa r^{1/2}}{\sqrt{c\left(m\wedge n\right)}}.$
This implies that under $A_{\delta}$, $L\leq\frac{1}{\omega_{\min}}\left(\mu^{2}r\frac{m+n-\mu^{2}r}{mn}+\frac{3C\mu\kappa r^{1/2}}{\sqrt{c\left(m\wedge n\right)}}\right)\leq C_{2}\frac{\mu^{2}\kappa r\left(m+n\right)}{\omega_{\min}mn}$ for some universal constant $C_{2}.$
@@
Choice of $T_{1}$. Note that the high probability bound of Lemma D.3 and D.5 together suggest that as long as $T_{1}=O\left(T^{\beta}\right)$ for $1/2<\beta\leq 1$, the OPE error that originates from the misspecification error will be of higher order than the first term with high probability. Furthermore, to ensure that the linear error term and the misspecification error term both have the best possible scaling in $T$, we should have $T_{1}=\Theta(T)$ and $T_{2}=\Theta(T).$ Specifically, for some $\alpha\in\left(0,1\right),$ we can choose $T_{1}=\left\lfloor\alpha T\right\rfloor$ and $T_{2}=\left\lceil\left(1-\alpha\right)T\right\rceil$.
@@
$\alpha T$ $\geq 1+\frac{c\mu^{4}\kappa^{2}r^{2}(m+n)}{\omega_{\min}mn}\log^{3}\left((m+n)\alpha T/\delta\right),$

### v2

@@ C.3 Additional Lemmas
  Input: Samples $\left(i_{t},j_{t},r_{t}\right)_{t\in[T]}$, context distribution $\rho$, behavior policy $\pi^{b}$, target policy $\pi$, sample size $T_{1}$, regularization parameter $\tau$
@@
  Output: $\hat{v}^{\pi}_{\text{\tt RS-PE}}=\sum_{i,j}\omega^{\pi}_{i,j}\phi_{i,j}^{T}\hat{\theta}$.
The general idea of the proof of Theorem 6.2 is to decompose the PE error of RS-PE as a sum of a term that depends on the misspecification error and a term incurred by the error between $\hat{\theta}$ and $\theta$. The analysis of the latter term, which is in fact the largest contributor to the overall error, is heavily inspired by the analysis in Appendix C.2. We first derive a high probability guarantee with the lower-order term that one would expect from the contextual linear bandit analysis, i.e., roughly $\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\sqrt{\frac{\log\left(1/\delta\right)}{T}}$. This bound is however not satisfactory in itself since it is a function of the randomness of the first phase of the algorithm. This issue can be resolved by leveraging the subspace recovery guarantees of Theorem 4.1 to show that $\|\phi_{\pi}\|_{2}$ converges towards $\|\psi_{\pi}\|_{2}$.
@@
Assume that $T_{1}=\lfloor\alpha T\rfloor$ for some $\alpha\in\left(0,1\right)$ and that the regularization parameter of $\widehat{\Lambda}_{\tau}$ verifies $\tau\leq\frac{r}{\left(m\wedge n\right)}\frac{\omega_{\min}}{\omega_{\max}}\log\left(16d/\delta\right)$. There exists a universal constant $C>0$ such that as long as $T\geq\frac{CL^{2}\kappa^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$ and $T\geq\frac{C}{\omega_{\min}}\log\left(16d/\delta\right)$, then with probability larger than $1-\delta$,
$\left|v^{\pi}-\hat{v}_{\text{\tt RS-PE}}\right|\leq\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\sqrt{\frac{2\sigma^{2}\log(16/\delta)}{\left(1-\alpha\right)T}}+K_{0}\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T}$
where
$K_{0}\lesssim\frac{L^{2}}{\|M\|_{\max}}\frac{\mu^{6}\kappa^{4}r^{3}\left(m+n\right)^{2}\sqrt{mn}}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}.$ (28)
@@
$\leq\left|\frac{1}{T_{2}}\sum_{t=T_{1}+1}^{T}\phi_{\pi}^{\top}\left(\xi_{t}+\epsilon_{i_{t},j_{t}}\right)\widehat{\Lambda}_{\tau}^{-1}\phi_{i_{t},j_{t}}\right|+\frac{\tau}{T_{2}}\phi_{\pi}^{\top}\widehat{\Lambda}_{\tau}^{-1}\theta+\epsilon_{\max}$
@@
$+\epsilon_{\max}\left(1+\left|\frac{1}{T_{2}}\sum_{t=T_{1}+1}^{T}\phi_{\pi}^{\top}\widehat{\Lambda}_{\tau}^{-1}\phi_{i_{t},j_{t}}\right|\right).$
@@
$L_{3}=\epsilon_{\max}\left(1+\left|\frac{1}{T_{2}}\sum_{t=T_{1}+1}^{T}\phi_{\pi}^{\top}\widehat{\Lambda}_{\tau}^{-1}\phi_{i_{t},j_{t}}\right|\right).$
As discussed before, we first perform the analysis conditionally on the first phase of the algorithm. We will then control the misspecification error induced by the first phase with Corollary 5.1. Let $L_{\phi}:=\max_{i,j}\phi_{i,j}^{\top}\Lambda_{\phi}^{-1}\phi_{i,j}$, $L_{\theta}:=\max_{i,j}|\phi_{i,j}^{\top}\theta|$. Adapting the proof of Theorem C.7 yields the following lemma.
@@
$|L_{3}|\leq\epsilon_{\max}\left(1+\sqrt{L_{\phi}}\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\left(1+4\sqrt{\frac{\log\left(16d/\delta\right)}{T_{2}}}\right)\right)$
with $E_{1}:=\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\sqrt{2\sigma^{2}\log(16/\delta)},E_{2}:=\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\left(L_{\phi}L_{\theta}+\sigma(1+4\sqrt{2d})\sqrt{L_{\phi}}\right)\log\left(32(d+1)/\delta\right),E_{3}:=\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\left(\frac{\sigma\sqrt{2}}{3}+4\sigma\sqrt{d}+4\sqrt{L_{\phi}}L_{\theta}\right)L_{\phi}\log^{3/2}\left(32(d+1)/\delta\right),E_{4}:=\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\frac{4\sigma\sqrt{2}}{3}\sqrt{d}L_{\phi}^{3/2}\log^{2}\left(32(d+1)/\delta\right).$
@@
$\left|L_{2}\right|$ $\leq\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\left(\|\Lambda_{\phi}^{1/2}\Delta\Lambda_{\phi}^{1/2}\|_{{\textup{op}}}\|\Lambda_{\phi}^{-1/2}W\|_{2}+\frac{\tau L_{\theta}}{\lambda_{\min}\left(\Lambda_{\phi}\right)T_{2}}\left(1+\|\Lambda^{1/2}\Delta\Lambda_{\phi}^{1/2}\|_{{\textup{op}}}\right)\right).$
for $W=\frac{1}{T_{2}}\sum_{t=T_{1}+1}^{T_{2}}\xi_{t}\phi_{i_{t},j_{t}}.$ The only difference is the additional term $L_{3}$, which can be controlled similarly to $L_{2}$. Indeed, by Cauchy-Schwartz and the definition of the operator norm,
@@
$\leq\frac{1}{T_{2}}\sum_{t=T_{1}+1}^{T}\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\sqrt{\phi_{i_{t},j_{t}}^{\top}\Lambda_{\phi}^{-1}\phi_{i_{t},j_{t}}}$
$+\frac{1}{T_{2}}\sum_{t=T_{1}+1}^{T}\|\Lambda_{\phi}^{-1/2}\phi_{\pi}\|_{2}\|\Lambda_{\phi}^{1/2}\Delta\Lambda_{\phi}^{1/2}\|_{{\textup{op}}}\|\Lambda_{\phi}^{-1/2}\phi_{i_{t},j_{t}}\|_{2}$
$\leq\sqrt{L_{\phi}}\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\left(1+\|\Lambda_{\phi}^{1/2}\Delta\Lambda_{\phi}^{1/2}\|_{{\textup{op}}}\right).$
Thus, $L_{3}\leq\epsilon_{\max}\left(1+\sqrt{L_{\phi}}\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\left(1+\|\Lambda_{\phi}^{1/2}\Delta\Lambda_{\phi}^{1/2}\|_{{\textup{op}}}\right)\right).$
By the same arguments as in the proof of Lemma B.6. from (Duan et al., 2020), if $T_{2}\geq 18L_{\phi}\log\left(16d/\delta\right)$ and $\tau\leq L_{\phi}\log(16d/\delta)\lambda_{\min}\left(\Lambda_{\phi}\right)$, then $\|\Lambda_{\phi}^{1/2}\Delta\Lambda_{\phi}^{1/2}\|_{{\textup{op}}}\leq 4\sqrt{\frac{\log(16d/\delta)L_{\phi}}{T_{2}}}$ with probability larger than $81-\delta/$ 8.
We can now derive high probability bounds on $L_{1}$, $L_{2}$ and $L_{3}$ by adapting the contextual linear bandit analysis. The main difference is that here, the analysis is performed conditionally on the randomness of the first phase of the algorithm. One subtlety to note is that we need to use a version of Proposition C.9 that is conditional on the first $T_{1}$ samples because the choices of $\gamma$ and $\varepsilon$ in the analysis depend on $\phi$, which depends on the estimator $\widehat{M}$ built with the first $T_{1}$ samples. It is not an issue as it directly implies an unconditional high probability bound by taking the expectation.
Following the proof of Theorem C.7 thus yields the high probability bound in the statement as soon as $T_{2}\geq 18L_{\phi}\log\left(16d/\delta\right)$ and $\tau\leq L_{\phi}\log(16d/\delta)\lambda_{\min}\left(\Lambda_{\phi}\right)$. All that remains to be checked is that these inequalities are true as soon as $T_{2}\geq\frac{18}{\omega_{\min}}\log\left(16d/\delta\right)$ and $\tau\leq\frac{r}{\left(m\wedge n\right)}\frac{\omega_{\min}}{\omega_{\max}}\log\left(16d/\delta\right)$. By Lemma D.11 and the fact that $\max_{i,j}\|\phi_{i,j}\|^{2}_{2}\geq\max\left(\|U\|^{2}_{2\to\infty},\|V\|^{2}_{2\to\infty}\right)$,
$\frac{r}{\left(m\wedge n\right)\omega_{\max}}\leq\max_{i,j}\frac{\|\phi_{i,j}\|^{2}_{2}}{\lambda_{\max}\left(\Lambda_{\phi}\right)}\leq L_{\phi}\leq\max_{i,j}\frac{\|\phi_{i,j}\|^{2}_{2}}{\lambda_{\min}\left(\Lambda_{\phi}\right)}\leq\frac{1}{\omega_{\min}},$
@@
Note that in addition to $L_{\phi}\leq\frac{1}{\omega_{\min}}$, we also have $L_{\theta}\leq\max_{i,j}\|\phi_{i,j}\|_{2}\|\theta\|_{2}\leq\sqrt{mn}\|M\|_{\max}$. This shows that the terms that depend on $E_{2}$, $E_{3}$ and $E_{4}$ are of lower order compared to the term that depends on $E_{1}$, and that they have a polynomial dependency in the model parameters (note that $\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\leq\sqrt{L_{\phi}}$). We will now see that we can obtain a sharper control of the dependency in the model parameters by leveraging high probability bounds on $L_{\phi}$ and $L_{\theta}$ instead of deterministic ones. These high probability bounds come freely with the control of the misspecification error.
@@
Let $\delta>0$. There exists an universal constant $c>0$ such that $T_{1}\geq\frac{cL^{2}\kappa^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right)$ ensures that with probability larger than $1-\delta$, the following inequalities happen at the same time :
  • $\epsilon_{\max}\lesssim\frac{L^{2}}{\sigma_{r}}\frac{\mu^{2}\kappa^{3}r(m+n)}{T_{1}\omega_{\min}(m\wedge n)}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right)$
  • $L_{\theta}\lesssim\frac{\sigma_{1}\mu^{2}r}{m\wedge n}$
  • $L_{\phi}\lesssim\frac{\mu^{2}r\left(m+n\right)}{\omega_{\min}mn}$.
@@
Since $T_{1}\gtrsim\frac{L^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right)$, by Theorem 4.1, there is a universal constant $C$ such that the event
@@
$A_{\delta}=\left\{\epsilon_{\textup{Sub-Rec}}\leq\frac{CL}{\sigma_{r}}\sqrt{\frac{\mu^{2}\kappa^{2}r(m+n)}{T_{1}\omega_{\min}(m\wedge n)}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right)}\right\}$
holds with probability at least $1-\delta.$ Under $A_{\delta}$, from (8),
$\epsilon_{\max}\leq\sigma_{1}\epsilon^{2}_{\textup{Sub-Rec}}\lesssim\frac{L^{2}}{\sigma_{r}}\frac{\mu^{2}\kappa^{3}r(m+n)}{T_{1}\omega_{\min}(m\wedge n)}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right).$
@@
Let us now prove that both of the remaining inequalities are true under $A_{\delta}.$ Specifically, we will see that $L_{\theta}$ (resp. $L_{\phi}$) can be deterministically bounded by a function of $\epsilon_{\max}$ (resp. $\epsilon_{\textup{Sub-Rec}}$).
@@
Control of $L_{\theta}$. Recall that $M_{i,j}=\phi_{i,j}^{\top}\theta+\epsilon_{i,j}.$ Consequently, $L_{\theta}=\displaystyle\max_{i,j}\left|\phi_{i,j}^{\top}\theta\right|\leq\|M\|_{\max}+\epsilon_{\max}$. Since $T_{1}\gtrsim\frac{L^{2}\kappa^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right)$, under $A_{\delta}$ we have $\epsilon_{\max}\lesssim\sigma_{1}\frac{\mu^{2}r}{m\wedge n}$, so that by Lemma B.8,
$L_{\theta}$ $\leq\|M\|_{\max}+\epsilon_{\max}$
$\lesssim\sigma_{1}\frac{\mu^{2}r}{\sqrt{mn}}+\sigma_{1}\frac{\mu^{2}r}{m\wedge n}$
$\lesssim\frac{\sigma_{1}\mu^{2}r}{m\wedge n}.$
@@
Control of $L_{\phi}$. Recall that $L_{\phi}\leq\displaystyle\max_{i,j}\frac{\|\phi_{i,j}\|^{2}_{2}}{\lambda_{\min}\left(\Lambda_{\phi}\right)}\leq\max_{i,j}\frac{\|\phi_{i,j}\|^{2}_{2}}{\omega_{\min}}.$ Note that by Lemma D.12, $\max_{i,j}\|\phi_{i,j}\|^{2}_{2}\leq\mu^{2}r\frac{m+n-\mu^{2}r}{mn}+3\epsilon_{\textup{Sub-Rec}}$. Consequently, $L_{\phi}\leq\frac{1}{\omega_{\min}}\left(\mu^{2}r\frac{m+n-\mu^{2}r}{mn}+3\epsilon_{\textup{Sub-Rec}}\right).$ Since $T_{1}\gtrsim\frac{L^{2}\kappa^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T_{1}}{\delta}\right)$, under $A_{\delta}$, we have $\epsilon_{\textup{Sub-Rec}}\lesssim\frac{\mu\sqrt{r}}{\sqrt{m\wedge n}},$ so
$L_{\phi}\lesssim\frac{1}{\omega_{\min}}\left(\mu^{2}r\frac{m+n-\mu^{2}r}{mn}+\frac{\mu\sqrt{r}}{\sqrt{m\wedge n}}\right)\lesssim\frac{\mu^{2}r\left(m+n\right)}{\omega_{\min}mn}.$
@@
Choice of $T_{1}$. Note that the high probability bounds of Lemma D.3 and D.5 together suggest that as long as $T_{1}=O\left(T^{\beta}\right)$ for $1/2<\beta\leq 1$, the OPE error that originates from the misspecification will be of higher order than the error induced by the least squares estimation of $\theta$ with high probability. Furthermore, to ensure that the least squares error term and the misspecification error term both have the best achievable scaling in $T$, we should have $T_{1}=\Theta(T)$ and $T_{2}=\Theta(T).$ Specifically, for some $\alpha\in\left(0,1\right),$ we can choose $T_{1}=\left\lfloor\alpha T\right\rfloor$ and $T_{2}=\left\lceil\left(1-\alpha\right)T\right\rceil$.
@@
$\alpha T$ $\geq 1+\frac{cL^{2}\kappa^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)\alpha T}{\delta}\right),$